OpenStudy (anonymous):

find the domain : g(x)= x+2 over x @Mathematics

OpenStudy (anonymous):

\[x \in \mathbb{R}\]

OpenStudy (anonymous):

umm no

OpenStudy (anonymous):

Yes. \[-\infty <x<\infty\]

OpenStudy (anonymous):

\[(-\infty,0) \cup (0, \infty)\]